# Time dependence of neutrino quantum kinetics in a core-collapse supernova

### v1

@@ Abstract
Our understanding of neutrino flavor conversion in the supernova core is still preliminary, despite its likely relevance to the neutrino-driven supernova mechanism. We present multi-angle and multi-energy numerical simulations of neutrino quantum kinetics within a spherically symmetric shell in the proximity of the region of neutrino decoupling. We rely on inputs from a one-dimensional core-collapse supernova model with a mass of and find that, at early post-bounce times ( s), no crossing is present in the angular distribution of the electron neutrino lepton number and flavor conversion develops due to vacuum mixing. Angular crossings appear for s and fast flavor conversion leads to flavor equipartition, with the spectral energy distribution of () and () becoming comparable. Notably, flavor equipartition is not a generic outcome of fast flavor conversion, but rather a consequence of the relatively similar properties of neutrinos of different flavors characterizing the late accretion phase. Artificially tweaking the collision term to introduce an electron lepton number angular crossing for s, we observe that flavor equipartition is not achieved. While our findings are restricted to a specific supernova model, they suggest a rich phenomenology of flavor conversion in the supernova core as a function of the post-bounce time which needs to be further explored to assess its impact on the explosion mechanism.
@@ 1 Introduction
The modification of neutrino flavor evolution due to refraction with other neutrinos can be broadly classified into slow [8, 9, 4], fast [10, 11, 12, 13, 14, 6, 7], and collisional [15]. Slow flavor evolution is differentiated from the fast one by the role of the vacuum frequency. In fact, fast flavor evolution can manifest in the limit of vanishing vacuum frequency and it is triggered by a crossing in the electron lepton number (ELN) angular distribution of neutrinos [14, 13, 16], whereas slow flavor evolution requires a nonzero vacuum term [9, 17]. Fast flavor evolution, as the name suggests, occurs over very small time scales—the characteristic frequency being directly proportional to the number density of neutrinos. On the other hand, collisional instabilities should be triggered by neutrino collisions with the medium in the high-density region in the proximity of neutrino decoupling [18, 19, 20, 21], but their relevance on flavor conversion physics within realistic astrophysical environments remains to be assessed [22, 23].
@@
Neutrinos play a crucial role in the core-collapse supernova mechanism [24, 25], as recently confirmed by multi-dimensional hydrodynamic simulations [26, 27, 28, 29]. Whether neutrino flavor conversion, triggered by any of the flavor instabilities highlighted above, affects the core collapse mechanism is the subject of ongoing research [30, 31, 32]. Preliminary work suggests that, if flavor equilibration should occur as a result of such instabilities, this would facilitate or hinder the supernova explosion according to the mass of the collapsing star [31]. Because of the technical challenges linked to the solution of the neutrino quantum kinetic equations (QKEs) and the vast difference in the characteristic scales typical of the source hydrodynamics and the ones entering the neutrino equations of motion, a self-consistent solution of this problem is not yet available [6, 7]. Nevertheless, it is possible to search for general trends by solving the QKEs within a simulation shell, taking into account the interplay among flavor evolution, collisions, and advection. In this case, a quasi-steady state configuration is found in the aftermath of flavor conversion [33, 34, 35, 32, 22, 36, 37, 38].
@@
An open question concerns whether flavor equipartition is a general solution of the QKEs. While recent work seems to point in this direction [39, 36, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49], the boundary conditions as well as the ELN distribution entering the solution of the QKEs could drastically affect the final flavor configuration [37, 33, 34]. In this paper, we investigate the neutrino flavor evolution in spherical symmetry, relying on input from several post-bounce time snapshots within s of a one-dimensional hydrodynamic simulation of a core-collapse supernova with a mass of [50]. We find that, in the early accretion phase ( s), flavor conversion is driven by vacuum mixing; while, in the late accretion phase ( s), equipartition between the neutrino spectral energy distributions of and as well as and . Yet, we confirm early findings [33, 34, 37] showing that flavor equipartition is not a general flavor outcome for any given unstable flavor configuration.
@@
This paper is organized as follows. Section 2 introduces the neutrino QKEs and outlines the approach adopted to model the collision term as well as the simulation setup. In Sec. 3, we describe the main features of our benchmark one-dimensional supernova model, compute the neutrino angular distributions of all flavors employing a multi-energy transport scheme and illustrate as they evolve as functions of the post-bounce time. Our main findings are presented in Sec. 4. Finally, we discuss our results in Sec. 5 and conclude in Sec. 6.
@@ 2.1 Quantum kinetic equations
The equations of motion that describe the evolution of the (anti)neutrino density matrices are:
@@
On the right hand side of Eqs. 2.1 and 2.2, the commutator of the Hamiltonian and the density matrix takes into account neutrino flavor conversion physics. The Hamiltonian consists of three components: the vacuum, matter, and self-interaction terms. The matter term effectively suppresses the mixing angle in the vacuum term; hence, we choose to adopt an effective vacuum mixing angle, ignoring the matter term:
@@
The vacuum frequency is with eV2, and rad denotes the matter suppressed mixing angle. The self-interaction Hamiltonian, unlike the vacuum Hamiltonian, depends on and is obtained integrating over all momentum modes. The integration over the azimuthal angle results in a constant factor that has been absorbed in the definition of the density matrices. On the other hand, the self-interaction Hamiltonian does not depend on energy. Following the attenuation method introduced in Ref. [51], we rescale the strength of neutrino self-interaction by a constant factor . The choice of is such that the self-interaction strength is always larger than at neutrino decoupling and the quasi-steady state flavor state configuration is likely not affected, but the time needed for reaching such a configuration changes with the computational time being reduced [52, 53].
The collision term on the right hand side of Eqs. 2.1 and 2.2, represented by and , includes three contributions: the emission term (), the absorption term (), and the direction changing term () [54, 55, 28]. For all flavors of neutrinos, pair production and Bremsstrahlung contribute to the emission and absorption terms, whereas the beta processes only contribute to electron-type neutrinos and antineutrinos. The direction changing term is given by the neutral current interaction between neutrinos and nucleons, which are assumed to be elastic. The inelastic contribution is negligible for the electron type neutrinos and thus does not affect the presence of ELN crossings. The collision terms are defined as follows:
@@ 2.2 Simulation setup
In order to solve the QKEs, we use the Julia implementation of the Adam-Moultan multistep integrator, with adaptive order and adaptive step size for the time variable 222The length scale characteristic of our neutrino ensemble is not given by the inverse of the self-interaction potential, rather it is a combination of the collision term and the self-interaction potential [33, 34]. Yet, it is important to ensure that the time step is much smaller than the inverse of the neutrino self-interaction strength; we do this by implementing an adaptive step size. (VCABM solver). The spatial derivatives in the advective term are calculated using the central difference method. To discretize the density matrix, we employ angle bins, energy bins for MeV, and radial bins for (with ans being the innermost and outermost radii of our simulation shell; see Fig. 1).
Our goal is to compute the “quasi-steady state” configuration achieved by the system 333A quasi-steady state configuration does not coincide with the solution of the boundary value problem obtained by setting the time derivative to zero. The quasi-steady state is also not approximately close to the solution of that boundary value problem. The reason for this counter-intuitive assertion is that, if Eqs. 2.1 and 2.2 are evolved for a long enough time, , and reach a steady state, but the phase of does not converge towards a steady state. This explains why we need to evolve the system as a function of time instead of attempting to find a solution to Eqs. 2.1 and 2.2 setting the time derivative to zero.. The latter should be independent of the initial configuration, but the choice of the initial configuration does affect the efficiency and numerical stability of the numerical solution. As pointed out in Refs. [33, 34], the most efficient way to tackle this issue numerically is to first solve Eqs. 2.1 and 2.2 for to obtain the “classical steady state” solution (i.e., we solve the classical Boltzmann equation). We then use the classical steady state solution as the initial configuration to solve Eqs. 2.1 and 2.2 for and . This approach has two advantages. First, it is easier to highlight the impact of flavor evolution; second, the initial classical steady state configuration is not too different from the final quasi-steady state. We stress, however, that the choice to use the classical solution as the initial configuration is not unique, in principle it might be possible to devise a better initial configuration.
@@ 3 Benchmark core-collapse supernova model
Since the neutrino angular distributions are not provided as an output of our benchmark hydrodynamic simulation, we compute them following the same procedure illustrated in Sec. III of Ref. [23], i.e. we set and as indicated in Table 1 for each snapshot and solve Eqs. 2.1 and 2.2 for . In contrast to Ref. [23], we compute the angular distributions of (anti)neutrinos solving Eqs. 2.1 and 2.2 for multiple energy modes (see Sec. 2.2).
The left panels of Figs. 3 and 4 display the contour plot of the energy-integrated in the plane spanned by and in the absence of flavor conversion (i.e., the classical steady state configuration). One can see that, at small radii (before neutrino decoupling), the angular distributions of neutrinos and antineutrinos are isotropic with an overall excess of . At larger radii, (anti)neutrinos start to decouple from matter entering the free streaming regime. As expected, because of the overall decrease of baryon density, neutrinos decouple at smaller radii as the post-bounce time increases. As (anti)neutrinos approach the free-streaming regime, an ELN crossing forms for s (visible from the white contour separating the red from the blue regions in the left panels of Fig. 4). The presence of an ELN crossing is a necessary condition for flavor instability and becomes a sufficient condition, if periodic boundary conditions are assumed [14, 59].
@@
4 Time dependent neutrino quantum kinetics
@@ 4.1 Early accretion phase: no ELN crossings, minimal flavor conversion
Although we do not find ELN crossings at early times ( s), this does not imply the absence of flavor evolution as displayed in the right panels of Fig. 3 for and s. Such flavor evolution is triggered by the vacuum mass term. Earlier work concluded that slow flavor evolution should be suppressed at early post-bounce times due to the large ratio of the self-interaction strength and the vacuum term [60, 61, 62, 63, 64, 65]. However, as recently shown in Ref. [66], if the ELN distribution is on the verge of developing a crossing, the vacuum term can induce a flavor instability even when the self-interaction strength is large (inducing a crossing in the effective ELN distribution). This is exactly what happens at and s. We verify this conjecture by performing the linear stability analysis without rescaling the self-interaction strength. As shown in Fig. 6, a non-zero growth rate of the off-diagonal term of the density matrix is found.
The right panels of Fig. 3 show that, unlike for the flavor instability induced by an ELN crossing (see Sec. 4.2), flavor equipartition does not take place. This can be attributed to a growth rate that does not dominate over the advective length scale as in the case of fast flavor instability and the different shape of the ELN angular distribution (which exhibits a larger difference between the properties of and with respect to later post-bounce times).
@@ 4.2 Late accretion phase: appearance of ELN crossings and fast flavor conversion
The right panels of Fig. 4 show the contour plots of the energy-integrated at the simulation time s in the plane spanned by and and in the presence of flavor conversion for , and s (from top to bottom respectively). We find that flavor conversion develops in the proximity of the ELN crossings (the latter being denoted by the white region in between the red and blue ones) for , and s. Flavor conversion is therefore driven by fast flavor instabilities, as expected from the preliminary findings of Ref. [23] in the single-energy approximation. In fact, Ref. [23] found that the growth rate of the fast flavor instability is four orders of magnitude larger than the collisional instability one for these time snapshots (cf. their Fig. 6). Reference [23] also found that collisional instabilities are present within small spatial regions at smaller radii, yet we do not see any significant flavor conversion linked to them.
In order to investigate whether flavor equipartition is achieved as a result of flavor conversion, Fig. 7 displays the energy-integrated for the quasi-steady-state solution obtained after flavor conversion for neutrinos (antineutrinos) on the left (right) panels. Note that, in the region of neutrino trapping, there is an excess of over due to the positive chemical potential of ; while the opposite is true for antineutrinos. For all analyzed configurations for s, we find flavor equipartition is achieved for neutrinos and antineutrinos. This is also visible from the left panels of Fig. 9, where the radial evolution of the neutrino number density of the different flavors after flavor conversion (quantum solution) is compared to the one obtained neglecting flavor conversion (classical solution). In particular, as shown in the right panels of Fig. 9 for (top two panels) and s (bottom two panels), the angle-integrated energy distribution of () is identical to the one of () in the quasi-steady state configuration and at large radii ( has been chosen to extract the spectral distributions plotted in Fig. 9, although any radius after neutrino decoupling shows a similar trend).

### v2

# Neutrino quantum kinetics in a core-collapse supernova

@@ Abstract
Our understanding of neutrino flavor conversion in the supernova core is still preliminary, despite its likely relevance to the neutrino-driven supernova mechanism. We present multi-angle and multi-energy numerical simulations of neutrino quantum kinetics within a spherically symmetric shell in the proximity of the region of neutrino decoupling. We rely on inputs from a one-dimensional core-collapse supernova model with a mass of and find that, at early post-bounce times ( s), no crossing is present in the angular distribution of the electron neutrino lepton number and flavor conversion is triggered by slow collective instabilities. Angular crossings appear for s and fast flavor conversion leads to flavor equipartition, with the spectral energy distribution of () and () becoming comparable. Notably, flavor equipartition is not a generic outcome of fast flavor conversion, rather it is a consequence of the relatively similar properties of neutrinos of different flavors characterizing the late accretion phase. Artificially tweaking the collision term to introduce an electron lepton number angular crossing for s, we observe that flavor equipartition is not achieved. While our findings are restricted to a specific supernova model, and they only take into account the feedback of the neutrino background on the flavor conversion, they suggest a rich phenomenology in the supernova core as a function of the post-bounce time which needs to be further explored to assess its impact on the explosion mechanism.
@@ 1 Introduction
The modification of neutrino flavor evolution due to refraction with other neutrinos can be broadly classified into slow [8, 9, 4], fast [10, 11, 12, 13, 14, 6, 7], and collisional [15]. Slow flavor evolution is differentiated from the fast one by the role of the vacuum frequency. In fact, fast flavor evolution can manifest in the limit of vanishing vacuum frequency and it is triggered by a crossing in the electron lepton number (ELN) angular distribution of neutrinos [14, 13, 16, 17], whereas slow flavor evolution requires a nonzero vacuum term [9, 18, 19]. Fast flavor evolution, as the name suggests, occurs over very small time scales—the characteristic frequency being directly proportional to the number density of neutrinos. On the other hand, collisional instabilities should be triggered by neutrino collisions with the medium in the high-density region in the proximity of neutrino decoupling [20, 21, 22, 23, 24], but their relevance on flavor conversion physics within realistic astrophysical environments remains to be assessed [25, 26].
@@
Neutrinos play a crucial role in the core-collapse supernova mechanism [27, 28], as recently confirmed by multi-dimensional hydrodynamic simulations [29, 30, 31, 32]. Whether neutrino flavor conversion, triggered by any of the flavor instabilities highlighted above, affects the core collapse mechanism is the subject of ongoing research [33, 34, 35]. Preliminary work suggests that, if flavor equilibration should occur as a result of such instabilities, this would facilitate or hinder the supernova explosion according to the mass of the collapsing star [34]. Because of the technical challenges linked to the solution of the neutrino quantum kinetic equations (QKEs) and the vast difference in the characteristic scales typical of the source hydrodynamics and the ones entering the neutrino equations of motion, a self-consistent solution of this problem is not yet available [6, 7, 36, 37]. Nevertheless, it is possible to search for general trends by solving the QKEs within a simulation shell, taking into account the interplay among flavor evolution, collisions, and advection. In this case, a quasi-steady state configuration is found in the aftermath of flavor conversion [38, 39, 40, 35, 25, 41, 42, 43].
@@
An open question concerns whether flavor equipartition (i.e. the angular and/or spectral distributions of the electron and non-electron flavors approach each other) is a general solution of the QKEs. While recent work seems to point in this direction [44, 41, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54], the boundary conditions as well as the ELN distribution entering the solution of the QKEs could drastically affect the final flavor configuration [42, 38, 39]. In this paper, we investigate the neutrino flavor evolution in spherical symmetry, relying on input from several post-bounce time snapshots within s of a one-dimensional hydrodynamic simulation of a core-collapse supernova with a mass of [55]. We find that, in the early accretion phase ( s), flavor conversion is triggered by slow collective effects; while, in the late accretion phase ( s), ELN crossings in the angular distribution lead to flavor conversion, with resulting equipartition between the neutrino spectral energy distributions of and as well as and . Yet, we confirm early findings [38, 39, 42] showing that flavor equipartition is not a general flavor outcome for any given unstable flavor configuration.
@@
This paper is organized as follows. Section 2 introduces the neutrino QKEs and outlines the approach adopted to model the collision term as well as the simulation setup. In Sec. 3, we describe the main features of our benchmark one-dimensional supernova model, compute the neutrino angular distributions of all flavors employing a multi-energy transport scheme and illustrate as they evolve as functions of the post-bounce time. Our main findings are presented in Sec. 4. Finally, we discuss our results in Sec. 5 and conclude in Sec. 6. Appendix A provides a comparison between the neutrino number densities computed in this work and the ones obtained from our benchmark hydrodynamical simulation.
@@ 2.1 Quantum kinetic equations
The equations of motion that describe the evolution of the (anti)neutrino density matrices are [56]:
@@
On the right hand side of Eqs. 2.1 and 2.2, the commutator of the Hamiltonian and the density matrix takes into account neutrino flavor conversion physics. The Hamiltonian consists of three components: the vacuum, matter, and self-interaction terms. For the sake of simplicity, we choose to neglect the matter term and adopt an effective mixing angle:
@@
The vacuum frequency is with eV2, and rad denotes the effective mixing angle. Although a small vacuum mixing angle is often adopted to mimic the effect of the matter background [18], whether flavor conversion is suppressed or not because of the large matter background [57, 58, 59, 60] should be subject of dedicated work for boundary problems like ours. In this work, we are only interested in exploring the flavor conversion outcome for different ELN configurations; however, we have tested that the quasi-steady state flavor configuration is unchanged for rad and rad (results not shown here). The self-interaction Hamiltonian, unlike the vacuum Hamiltonian, depends on and is obtained integrating over all momentum modes. The integration over the azimuthal angle results in a constant factor that has been absorbed in the definition of the density matrices. On the other hand, the self-interaction Hamiltonian does not depend on energy. Following the attenuation method introduced in Ref. [61], we rescale the strength of neutrino self-interaction by a constant factor . The choice of is such that the self-interaction strength is always larger than at neutrino decoupling and the quasi-steady state flavor state configuration is likely not affected, but the time needed for reaching such a configuration changes with the computational time being reduced [43, 62].
The collision term on the right hand side of Eqs. 2.1 and 2.2, represented by and , includes three contributions: the emission term (), the absorption term (), and the direction changing term () [63, 56, 64, 31]. For all flavors of neutrinos, pair production and Bremsstrahlung contribute to the emission and absorption terms, whereas the beta processes only contribute to electron-type neutrinos and antineutrinos. The direction changing term is given by the neutral current interaction between neutrinos and nucleons, which are assumed to be elastic. The inelastic contribution is negligible for the electron type neutrinos and thus does not affect the presence of ELN crossings. The collision terms are defined as follows:
@@ 2.2 Simulation setup
In order to solve the QKEs, we use the Julia implementation of the Adam-Moultan multistep integrator, with adaptive order and adaptive step size for the time variable 222The length scale characteristic of our neutrino ensemble is not given by the inverse of the self-interaction potential, rather it is a combination of the collision term and the self-interaction potential. We refer the interested reader to Sec. VI and Appendix A of Ref. [38] and Sec. IV and Appendix A of Ref. [39]. Yet, it is important to ensure that the time step is much smaller than the inverse of the neutrino self-interaction strength; we do this by implementing an adaptive step size. (VCABM solver). The spatial derivatives in the advective term are calculated using the central difference method. Numerical convergence (test results not shown here) has been obtained by employing angle bins, energy bins for MeV, and radial bins for (with ans being the innermost and outermost radii of our simulation shell; see Fig. 1).
Our goal is to compute the “quasi-steady state” configuration achieved by the system 333A quasi-steady state configuration does not coincide with the solution of the boundary value problem obtained by setting the time derivative to zero. The quasi-steady state is also not approximately close to the solution of that boundary value problem. The reason for this counter-intuitive assertion is that, if Eqs. 2.1 and 2.2 are evolved for a long enough time, , and reach a steady state, but the phase of does not converge towards a steady state. This explains why we need to evolve the system as a function of time instead of attempting to find a solution to Eqs. 2.1 and 2.2 setting the time derivative to zero.. The latter should be independent of the initial configuration (cf. Ref. [42]), but the choice of the initial configuration does affect the efficiency of the numerical solution. Future work should be devoted to assess the putative existence of such unique quasi-steady state configuration, see e.g. Ref. [67]. As pointed out in Refs. [38, 39], the most efficient way to tackle this issue numerically is to first solve Eqs. 2.1 and 2.2 for to obtain the “classical steady state” solution (i.e., we solve the classical Boltzmann equation). We then use the classical steady state solution as the initial configuration to solve Eqs. 2.1 and 2.2 for and . This approach has two advantages. First, it is easier to highlight the impact of flavor evolution; second, the initial classical steady state configuration is not too different from the final quasi-steady state. We stress, however, that the choice to use the classical solution as the initial configuration is not unique, in principle it might be possible to devise a better initial configuration.
@@ 3 Benchmark core-collapse supernova model
Since the neutrino angular distributions are not provided as an output of our benchmark hydrodynamic simulation, we compute them following the same procedure illustrated in Sec. III of Ref. [26], i.e. we set and as indicated in Table 1 for each snapshot and solve Eqs. 2.1 and 2.2 for . In contrast to Ref. [26], we compute the angular distributions of (anti)neutrinos solving Eqs. 2.1 and 2.2 for multiple energy modes (see Sec. 2.2). Appendix A provides a comparison between the neutrino number densities computed through this approach and the ones obtained from the supernova hydrodynamic simulation.
The left panels of Figs. 3 and 4 display the heatmaps of the energy-integrated in the plane spanned by and in the absence of flavor conversion (i.e., the classical steady state configuration). One can see that, at small radii (before neutrino decoupling), the angular distributions of neutrinos and antineutrinos are isotropic with an overall excess of . At larger radii, (anti)neutrinos start to decouple from matter entering the free streaming regime. As expected, because of the overall decrease of baryon density, neutrinos decouple at smaller radii as the post-bounce time increases. As (anti)neutrinos approach the free-streaming regime, an ELN crossing forms for s (visible from the white contour separating the red from the blue regions in the left panels of Fig. 4). The presence of an ELN crossing is a necessary condition for flavor instability and becomes a sufficient condition, if periodic boundary conditions are assumed [14, 69, 17].
@@
4 Neutrino quantum kinetics for different post-bounce times
@@ 4.1 Early accretion phase: no ELN crossings, minimal flavor conversion
Although we do not find ELN crossings at early times ( s), this does not imply the absence of flavor evolution as displayed in the right panels of Fig. 3 for and s. Such flavor evolution is triggered by the vacuum mass term. Earlier work concluded that slow flavor evolution may be suppressed at early post-bounce times due to the large ratio of the self-interaction strength and the vacuum term [57, 70, 71, 72, 73, 74]. However, as recently shown in Ref. [75], if the ELN distribution is on the verge of developing a crossing, the non-zero vacuum term and the neutrino-neutrino interaction term induce a collective instability. This is exactly what happens at and s. We verify this conjecture by performing the linear stability analysis without rescaling the self-interaction strength. As shown in Fig. 6, a non-zero growth rate of the off-diagonal term of the density matrix is found.
The right panels of Fig. 3 show that, unlike for the flavor instability induced by an ELN crossing (see Sec. 4.2), flavor equipartition does not take place. This can be attributed to a combination of the ELN configurations and the fact that substantial flavor conversion cannot take place before (anti)neutrinos are advected out of the simulation shell.
@@ 4.2 Late accretion phase: appearance of ELN crossings and fast flavor conversion
The right panels of Fig. 4 show the heatmaps of the energy-integrated at the simulation time s in the plane spanned by and and in the presence of flavor conversion for , and s (from top to bottom respectively). We find that flavor conversion develops in the proximity of the ELN crossings (the latter being denoted by the white region in between the red and blue ones) for , and s. Flavor conversion is therefore driven by fast flavor instabilities, as expected from the preliminary findings of Ref. [26] in the single-energy approximation. In fact, Ref. [26] found that the growth rate of the fast flavor instability is four orders of magnitude larger than the collisional instability one for these time snapshots (cf. their Fig. 6). Reference [26] also found that collisional instabilities are present within small spatial regions at smaller radii, yet we do not see any significant flavor conversion linked to them.
In order to investigate whether flavor equipartition is achieved as a result of flavor conversion 444 It should be noted that, in the presence of collisions, the lepton number is not conserved in the neutrino sector., Fig. 7 displays the energy-integrated for the quasi-steady-state solution obtained after flavor conversion for neutrinos (antineutrinos) on the left (right) panels. Note that, in the region of neutrino trapping, there is an excess of over due to the positive chemical potential of ; while the opposite is true for antineutrinos. For all analyzed configurations for s, we find flavor equipartition (i.e. the angular and spectral energy distributions of the and flavors approach each other, becoming comparable) is achieved for neutrinos and antineutrinos. This is also visible from the left panels of Fig. 9, where the radial evolution of the neutrino number density of the different flavors after flavor conversion (quantum solution) is compared to the one obtained neglecting flavor conversion (classical solution). In particular, as shown in the right panels of Fig. 9 for (top two panels) and s (bottom two panels), the angle-integrated energy distribution of () is identical to the one of () in the quasi-steady state configuration and at large radii ( has been chosen to extract the spectral distributions plotted in Fig. 9, although any radius after neutrino decoupling shows a similar trend).